OpenStudy (anonymous):

The depth (inches) of water that accumulates in the spring soil from melted snow can be determined by dividing the cumulative winter snowfall (inches) by 12. Determine the amount of winter snow that accumulates 6 inches of water in the soil.

OpenStudy (anonymous):

seriously though, not 12th grade

OpenStudy (anonymous):

6/12

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

misread

OpenStudy (anonymous):

6 * 12, easier than originally expected

OpenStudy (anonymous):

72

OpenStudy (anonymous):

Bill bought a new cellular phone. His monthly bill will be $32.99 each month which includes 100 minutes of use. Additional minutes will cost $0.19 per minute. Write a linear model that represents the amount of the bill, C, based on the total number of minutes, m, the phone is used in a month. Assume m 4072-01-02-03-00_files/i0060000.jpg 100 minutes. a. C = 0.19m + 32.99 c. C = 0.19m b. C = 32.99(m - 100) + 0.19 d. C = 0.19(m - 100) + 32.99 thats the last one i believe